# Can It Edit? Evaluating the Ability of Large Language Models to Follow Code Editing Instructions

### 29 Dec 2023 v3

@@ Abstract.
We introduce a carefully crafted benchmark of code editing tasks and use it evaluate several cutting edge LLMs. Our evaluation exposes a significant gap between the capabilities of state-of-the-art open and closed models. For example, even GPT-3.5-Turbo is 8.8% better than the best open model at editing code.
@@ 1. Introduction
Large language models of code (Code LLMs) are starting to become an essential tool for software engineering practice and research. There has been significant research on synthesizing code from natural language instructions, but comparatively less attention has been given to code editing tasks. However, LLM users expect models to be capable of editing code. For example, the LMsys dataset of real-world conversations with chatbots (Zheng et al., 2023) has 4,188 conversations with code, and 831 (19%) of these involve edits, where the user prompts the model to update generated code based on natural language instructions (Appendix C). In general, code editing encompasses activities like feature addition or removal, bug fixing, and code refactoring (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023).
@@
In this work, we introduce CanItEdit, a novel dataset comprising 54 hand-crafted instructional code editing problems. These problems, featuring both descriptive and lazy instructions, are coupled with an extensive hidden test suite. Designed to assess a model’s proficiency in handling realistic code editing scenarios, CanItEdit serves as a platform for evaluating state-of-the-art Code LLMs in instructional code editing. Our evaluation focuses on measuring the accuracy of a given model’s ability to write correct code modifications without introducing superfluous code. We conduct comprehensive assessments of closed and open models, revealing significant performance disparities between the leading closed and open models in this domain (Section 5). To help address this gap, we propose a training dataset and methodology for code editing. Our findings demonstrate that fine-tuning open Code LLMs on this dataset can significantly enhance their performance (Section 4).
@@
    We propose a novel metric, ExcessCode, for assessing code editing models. This metric quantifies the volume of unnecessary code produced by a model when generating a correct solution (Section 5.1).
@@
    We present a specially tailored dataset for code editing, along with an effective training methodology, demonstrating significantly enhanced code editing performance through fine-tuning a state-of-the-art Code LLM (Section 4).
The benchmark, models, training dataset, and the code to reproduce our work are available at:
https://github.com/nuprl/CanItEdit/
@@ 2. Related Work
Several benchmarks exist that test a model’s code generation ability. HumanEval and MBPP are two prominent benchmarks for evaluating Code LLMs in Python programming (Chen et al., 2021; Austin et al., 2021). MultiPL-E expands these benchmarks to 18+ additional programming languages (Cassano et al., 2023). These benchmarks assess model-generated candidate completions against a series of human-authored unit tests. EvalPlus (Liu et al., 2023) utilizes mutation testing to expand the test suites of the Python benchmarks. All of these benchmarks utilize the pass@k metric, which measures the likelihood of the model generating a completion that passes all of the tests in $k$ tries; we also adopt this metric in our evaluation (Section 5.1). However, these benchmarks are limited to the evaluation of a model’s ability to generate a single function from a natural language description and do not assess code editing capabilities. HumanEvalPack (Muennighoff et al., 2023) is a comprehensive benchmark designed for evaluating Code LLMs across various code generation tasks, such as synthesis, explanation for code understanding, and bug fixing. Specifically, HumanEvalFix, a bug-fixing variant of HumanEvalPack, is extensively used for assessing the models’ capabilities in code refinement (Moon et al., 2023; Muennighoff et al., 2023).
SWE-Bench (Jimenez et al., 2023) evaluates Code LLMs on a broad spectrum of tasks that are performed by real-world software engineers, and require planning, retrieval, code editing, and more for successful task completion. Our work is more narrowly focused on code editing, and we believe this focus will help guide model development. Another difference with SWE-Bench is that our benchmark is handcrafted, whereas SWE-Bench is based on PRs and issues from popular GitHub repositories. This increases the risk of contamination, particularly with models such as StarCoder, which is trained on several GBs of GitHub issues (Li et al., 2023a).
@@
Previous studies on code editing with large language models (LLMs) have predominantly focused on bug fixing (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023; Joshi et al., 2023; Wei et al., 2023), a specific subset of code editing, fill-in-the-middle code completion (Bavarian et al., 2022; Fried et al., 2023; Yee and Guha, 2023; Rozière et al., 2023; AI, 2023), an inference strategy that requires specific insert locations, and intrinsic code editing (Li et al., 2023b; Gupta et al., 2023), which involves editing code without a specified instruction, exerting the model’s ability to intrinsically ascertain the desired code changes. Recently, LLMs have progressed in code editing guided by natural language without specific edit locations (Hu et al., 2023; Li et al., 2023a; Muennighoff et al., 2023). However, this advancement lacks benchmark evaluations to effectively measure the models’ code editing skills. Notably, StarCoder (Li et al., 2023a), the first LLM trained on an extensive dataset of commits using the format <before><commit message><after>, has shown enhanced code editing capabilities (Section 5). Before this study, StarCoder’s practical code editing performance had not been assessed. The recent introduction of InstructCoder (Hu et al., 2023), a model explicitly trained and evaluated for code editing, marks a significant step towards code editing with LLMs. However, several concerns persist: 1) Its evaluation involved GPT-4 (OpenAI, 2023a) and human-provided labels on an unreleased dataset, which raises issues regarding reproducibility and comparability in future research; 2) Their approach of training on a dataset derived from ChatGPT interactions may potentially conflict with OpenAI’s usage terms (OpenAI, 2023c); 3) The model has not been publicly released, prohibiting us from evaluating it on our benchmark; 4) Our analysis of the training data, which was publicly released, reveals that the mean number of lines in the ‘before‘ and ‘after‘ code segments is $18.6$ with a standard deviation of $10.1$. This is not reflective of typical real-world code editing scenarios.
@@ 3. The CanItEdit Dataset
The function hello_world should return the string parameter ”name” converted to uppercase concatenated to the string ’ says, ”Hello World!”’. For example, hello_world(’the cow’) should return ’THE COW says, ”Hello World!”’. For another example, hello_world(’joe’) should return ’JOE says, ”Hello World!”’.
@@ 3.1. Problem Construction
CanItEdit features 54 Python code editing problems, each comprising a ‘before‘ and an ‘after‘ code segment, two types of natural language instructions (descriptive and lazy), and a hidden test suite. The task for models is to transform the ‘before‘ code segment into the ‘after‘ segment based on either instruction, aiming to pass the hidden tests. Inspired by HumanEval’s methodology (Austin et al., 2021), we hand-wrote the problems, avoiding public sources like GitHub to reduce pre-training exposure. We also verified that the instructions are unique to this dataset and not part of our fine-tuning data (section 4). Problems range from simple function edits to complex, multi-class challenges, covering data structures, algorithms, mathematics, language processing, and game programming. Some require popular external Python libraries like NumPy, Pandas, PyTorch, and Z3. Dataset statistics and example problems are detailed in Table 1 and Appendix B, respectively.
The ‘before‘ code segments in CanItEdit represent various starting states, ranging from functional programs needing additional features to those with bugs or incomplete implementations requiring fixes or optimizations. These segments are designed to mirror diverse real-world coding scenarios. Conversely, the ‘after‘ segments illustrate the correct solutions that fulfill the task requirements and clear the test suite.
@@
In both, the model must generate code that meets the instruction and passes the hidden tests. Descriptive instructions offer detailed guidance, including function names and input-output examples, while lazy instructions provide minimal information, requiring the model to infer user intent and rely more on the ‘before‘ segment. Both instructions should lead to an equivalent ‘after‘ segment. For instance, in the hello_world problem (Figure 2), the descriptive instruction is comprehensive, whereas the lazy instruction is brief, pushing the model to deduce user intent. Further discussions on human-written and Reflexion-generated instructions are in Appendix C.
@@ 3.2. Test Suites
    Completeness: Each suite comprehensively covers all inputs and edge cases. This includes numerous test cases per problem, targeting edge and corner cases, with $100\%$ code coverage verified using Coverage.py (Batchelder and Contributors to Coverage.py, [n. d.]). It is worth noting that code coverage is not as robust as mutation testing, which is employed by EvalPlus (Liu et al., 2023).
@@
    Correctness: The suites are bug-free, passing all tests with the ‘after‘ code while failing at least one with the ‘before‘ code.
@@
We hand-crafted test suites for each problem, incorporating diverse testing methods ranging from simple unit tests to complex property-based testing, mocking, fuzzing, and integration tests. For instance, one of our benchmark problems involves implementing a strategy for a Tic-Tac-Toe game that outperforms a baseline strategy (Figure 9). The lazy instruction for this problem is: Create a strategy ‘GoodStrategy‘, that beats ‘CornerStrategy‘. Do not modify the ‘Game‘ class. To test this, the suite includes unit tests for both the ‘Game‘ and ‘CornerStrategy‘ classes, along with integration tests that evaluate the entire program, ensuring that ‘GoodStrategy‘ wins over ‘CornerStrategy‘. Additionally, we use Python’s inspect module to check if the ‘Game‘ class remains unmodified, adhering to the problem’s constraints.
@@ 4. Fine-tuning
This section outlines our methodology for fine-tuning a Code LLM specifically for code editing tasks. We fine-tune a model based on the DeepSeek-Coder-Base family of Code LLMs (AI, 2023), which are variants of Code Llama (Rozière et al., 2023) trained from scratch on 2T tokens comprised of 87% permissively licensed code from GitHub and 13% natural language, using the same filtering rules as StarCoder’s data collection (Li et al., 2023a). At the time of writing, these models are the top-performing, open-access foundational Code LLMs, excelling in various code generation benchmarks. Furthermore, they are distributed under a permissive open-source license, allowing free use and modification for research and commercial purposes. We selected these base models because they exhibit robust performance on CanItEdit, even without being specifically trained for this or any other instructional tasks, thus highlighting their exceptional generalization capabilities (Section 5). We hand-crafted a training dataset for code editing, which we describe in Section 4.1, and fine-tuned the DeepSeek-Coder-Base 6.7 billion parameter model on this dataset, which we refer to as EditCoder.
@@ 4.1. Training Data
Figure 3. Sunburst plot of the top 20 most frequent initial verbs, with their corresponding top 10 root nouns, in the commit messages of EditPackFT.
EditPackFT Dataset Statistics
Total Commits 22,602
Unique Initial Verbs 184
Total Tokens 12M
Code Segments Mean Std. Dev.
Mean Lines 29.2 13.7
Commit Message Mean Std. Dev.
Mean Tokens 10.1 4.6
Table 2. Training dataset statistics for EditPackFT
Figure 4. The distribution of the number of lines in the ‘before‘ and ‘after‘ code segments in the EditPackFT dataset. The 99th percentile is removed for clarity.
We experiment with a training dataset for code editing, which we refer to as EditPackFT. We created the EditPackFT dataset by further filtering the Python split of the CommitPackFT dataset (Muennighoff et al., 2023), which was used to train OctoCoder.
CommitPack is an extensive dataset comprising 4TB of permissively licensed commits from a 2016 GitHub snapshot across various programming languages. CommitPackFT is a subset of CommitPack, filtered for it to be amenable to instruction-tune Code LLMs. The primary criterion for CommitPackFT’s selection involved retaining commits whose messages begin with an imperative verb, mirroring the typical structure of natural language instructions. We apply a series of additional filtering steps, which make the dataset more suitable for code editing. We remove any item that passes any of the following predicates:
@@
    The presence of an empty ‘before‘ or ‘after‘ code segment, disregarding whitespace.
@@
    The inclusion of the words TODO, FIXME, or BUG in the ‘after‘ code segment, which signals an incomplete commit.
@@
    A Levenshtein distance below $10$ between the ‘before‘ and ‘after‘ code segments, indicative of trivial changes.
@@
    Incorrect parsing of the ‘after‘ code using the Python ast module.
Originally, the dataset contained 56,025 commits, and after applying the filtering steps, we are left with 22,602 commits. As shown by Figure 4 and Table 2, the mean number of lines in the ‘before‘ and ‘after‘ code segments is $29.2$. We further analyzed the original CommitPackFT dataset, ensure that our filtering wasn’t the cause of the short code segments, and find that the mean number of lines is similar, with a mean of $28.8$. We recognize that while this distribution may be suitable for small-scale code editing tasks, it is not representative of real-world scenarios, where the code segments are typically longer and more complex. We also analyze the distribution of the commit message lengths, and find that the mean token count is $10.1$. Figure 3 illustrates a sunburst plot of the most frequent initial verbs in the commit messages of Commits2023FT, along with their corresponding root nouns. The set of initial verbs in EditPackFT is composed of $184$ unique verbs, and the most frequent verb is add, which appears in $6,278$ commits.

### 05 Mar 2024 v4

@@ Abstract.
We introduce a carefully crafted benchmark of code editing tasks and use it to evaluate several cutting edge LLMs. Our evaluation exposes a significant gap between the capabilities of state-of-the-art open and closed models. For example, even GPT-3.5-Turbo is 8.8% better than the best open model at editing code.
@@ 1. Introduction
Large language models of code (Code LLMs) are starting to become an essential tool for software engineering practice and research. There has been significant research on synthesizing code from natural language instructions, but comparatively less attention has been given to code editing tasks. However, LLM users expect models to be capable of editing code. For example, the LMsys dataset of in-the-wild conversations with chatbots (Zheng et al., 2023) has 4,188 conversations with code, and 831 (19%) of these involve edits, where the user prompts the model to update generated code based on natural language instructions (Appendix D). In general, code editing encompasses activities like feature addition or removal, bug fixing, and code refactoring (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023).
@@
In this work, we introduce CanItEdit, a novel dataset comprising 54 hand-crafted instructional code editing problems. These problems, featuring both descriptive and lazy instructions, are coupled with an extensive hidden test suite. Designed to assess a model’s proficiency in handling diverse code editing scenarios, CanItEdit serves as a platform for evaluating state-of-the-art Code LLMs in instructional code editing. Our evaluation focuses on measuring the accuracy of a given model’s ability to write correct code modifications without introducing superfluous code. We conduct comprehensive assessments of closed and open models, revealing significant performance disparities between the leading closed and open models in this domain (Section 5). To help address this gap, we propose a training dataset and methodology for code editing. Our findings demonstrate that fine-tuning open Code LLMs on this dataset can significantly enhance their performance (Section 4).
@@
    We propose a novel metric, ExcessCode, for assessing code editing models. This metric quantifies the volume of unused code produced by a model when generating a correct solution (Section 5.1).
@@
    We present a specially tailored dataset for code editing, along with an effective training methodology, demonstrating significantly enhanced code editing performance through fine-tuning models of three varying sizes (Section 4).
@@ 2. Related Work
Several benchmarks exist that test a model’s code generation ability. HumanEval and MBPP are two prominent benchmarks for evaluating Code LLMs in Python programming (Chen et al., 2021; Austin et al., 2021). MultiPL-E expands these benchmarks to 18+ additional programming languages (Cassano et al., 2023b). These benchmarks assess model-generated candidate completions against a series of human-authored unit tests. EvalPlus (Liu et al., 2023) utilizes mutation testing to expand the test suites of the Python benchmarks. All of these benchmarks utilize the pass@k metric, which measures the likelihood of the model generating a completion that passes all of the tests in $k$ tries; we also adopt this metric in our evaluation (Section 5.1). However, these benchmarks are limited to the evaluation of a model’s ability to generate a single function from a natural language description and do not assess code editing capabilities. HumanEvalPack (Muennighoff et al., 2023) is a comprehensive benchmark designed for evaluating Code LLMs across various code generation tasks, such as synthesis, explanation for code understanding, and bug fixing. Specifically, HumanEvalFix, a bug-fixing variant of HumanEvalPack, is extensively used for assessing the models’ capabilities in code refinement (Moon et al., 2023; Muennighoff et al., 2023).
SWE-Bench (Jimenez et al., 2023) evaluates Code LLMs on a broad spectrum of tasks that are performed in the wild by software engineers, and require planning, retrieval, code editing, and more for successful task completion. Our work is more narrowly focused on code editing, and we believe this focus will help guide model development. Another difference with SWE-Bench is that our benchmark is handcrafted, whereas SWE-Bench is based on PRs and issues from popular GitHub repositories. This increases the risk of contamination, particularly with models such as StarCoder, which is trained on several GBs of GitHub issues (Li et al., 2023a).
@@
Previous studies on code editing with large language models (LLMs) have predominantly focused on bug fixing (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023; Joshi et al., 2023; Wei et al., 2023), a specific subset of code editing, fill-in-the-middle code completion (Bavarian et al., 2022; Fried et al., 2023; Yee and Guha, 2023; Roziere et al., 2023; Guo et al., 2024), an inference strategy that requires specific insert locations, and intrinsic code editing (Li et al., 2023b; Gupta et al., 2023), which involves editing code without a specified instruction, exerting the model’s ability to intrinsically ascertain the desired code changes. Recently, LLMs have progressed in code editing guided by natural language without specific edit locations (Hu et al., 2023; Li et al., 2023a; Muennighoff et al., 2023). However, this advancement lacks benchmark evaluations to effectively measure the models’ code editing skills. Notably, StarCoder (Li et al., 2023a), the first LLM trained on an extensive dataset of commits using the format <before><commit message><after>, has shown enhanced code editing capabilities (Section 5). Before this study, StarCoder’s instructional code editing performance had not been evaluated. The recent introduction of InstructCoder (Hu et al., 2023), a model explicitly trained and evaluated for code editing, marks a significant step towards code editing with LLMs. However, its evaluation involved GPT-4 (OpenAI, 2023a) and human-provided labels on an unreleased dataset, which raises issues regarding reproducibility and comparability in future research and the model has not been publicly released, prohibiting us from evaluating it on our benchmark.
@@ 3. The CanItEdit Dataset
The function hello_world should return the string parameter ”name” converted to uppercase concatenated to the string ‘ says, ”Hello World!”’. For example, hello_world(‘the cow’) should return ‘THE COW says, ”Hello World!”’. For another example, hello_world(‘joe’) should return ‘JOE says, ”Hello World!”’.
@@ 3.1. Problem Construction
CanItEdit features 54 Python code editing problems, each comprising a ‘before’ and an ‘after’ code segment, two types of natural language instructions (descriptive and lazy), and a hidden test suite. The task for models is to transform the ‘before’ code segment into the ‘after’ segment based on either instruction, aiming to pass the hidden tests. Inspired by HumanEval’s methodology (Austin et al., 2021), we hand-wrote the problems, avoiding public sources like GitHub to reduce pre-training exposure. We also verified that the instructions are unique to this dataset and not part of our fine-tuning data (section 4). Problems range from simple function edits to complex, multi-class challenges, covering data structures, algorithms, mathematics, language processing, and game programming. Some require popular external Python libraries like NumPy, Pandas, PyTorch, and Z3. Dataset statistics and example problems are detailed in Table 1 and Appendix B, respectively.
The ‘before’ code segments in CanItEdit represent various starting states, ranging from functional programs needing additional features to those with bugs or incomplete implementations requiring fixes or optimizations. Conversely, the ‘after’ segments illustrate the correct solutions that fulfill the task requirements and clear the test suite.
@@
In both, the model must generate code that meets the instruction and passes the hidden tests. Descriptive instructions offer detailed guidance, including function names and input-output examples, while lazy instructions provide minimal information, requiring the model to infer user intent and rely more on the ‘before’ segment. Both instructions should lead to an equivalent ‘after’ segment. For instance, in the hello_world problem (Figure 2), the descriptive instruction is comprehensive, whereas the lazy instruction is brief, pushing the model to deduce user intent. Further discussions on human-written and Reflexion-generated instructions are in Appendix D.
@@ 3.2. Test Suites
    Completeness: Each suite comprehensively covers many inputs and edge cases. This includes numerous test cases per problem, targeting edge and corner cases, with $100\%$ code coverage verified using Coverage.py (Batchelder and Contributors to Coverage.py, [n. d.]). It is worth noting that code coverage is not as robust as mutation testing, which is employed by EvalPlus (Liu et al., 2023).
@@
    Correctness: The suites are bug-free, passing all tests with the ‘after’ code while failing at least one with the ‘before’ code.
@@
We hand-crafted test suites for each problem, incorporating diverse testing methods ranging from simple unit tests to complex property-based testing, mocking, fuzzing, and integration tests. For instance, one of our benchmark problems involves implementing a strategy for a Tic-Tac-Toe game that outperforms a baseline strategy (Figure 9). The lazy instruction for this problem is: Create a strategy ‘GoodStrategy’, that beats ‘CornerStrategy’. Do not modify the ‘Game’ class. To test this, the suite includes unit tests for both the ‘Game’ and ‘CornerStrategy’ classes, along with integration tests that evaluate the entire program, ensuring that ‘GoodStrategy’ wins over ‘CornerStrategy’. Additionally, we use Python’s inspect module to check if the ‘Game’ class remains unmodified, adhering to the problem’s constraints.
@@ 4. Fine-tuning
This section outlines our methodology for fine-tuning a Code LLM specifically for code editing tasks. We fine-tune models based on the DeepSeek-Coder-Base family of Code LLMs (Guo et al., 2024), which are variants of CodeLlama (Roziere et al., 2023) trained from scratch on 2T tokens comprised of 87% permissively licensed code from GitHub and 13% natural language, using the same filtering rules as StarCoder’s data collection (Li et al., 2023a). At the time of writing, these models are the top-performing, open-access foundational Code LLMs, excelling in various code generation benchmarks. Furthermore, they are distributed under a permissive open-source license, allowing free use and modification for research and commercial purposes. We selected these base models because they exhibit robust performance on CanItEdit, even without being specifically trained for this or any other instructional tasks, thus highlighting their exceptional generalization capabilities (Section 5).
For our ablation studies, we focus on the variant with 6.7 billion parameters. This model offers an ideal balance between size and performance, allowing us to extrapolate results to larger models with more parameters without the need for extensive training, which would be resource-intensive and time-consuming. Following the most efficient training strategy identified, we also fine-tune the 1.3b and 33b models to evaluate the impact of model size on code editing performance. Our fine-tuned models are referred to as EditCoder.
@@ 4.1. Training Data
Figure 3. Sunburst plot of the top 20 most frequent initial verbs, with their corresponding top 10 root nouns, in the commit messages of Commits2023FT.
(a) EditPackFT
(b) Commits2023FT
Figure 4. The distribution of the number of lines in the ‘before’ and ‘after’ code segments in the EditPackFT and Commits2023FT datasets. The 99th percentile is removed for clarity.
Dataset Statistics
EditPackFT Commits2023FT
Total Commits 22,602 24,129
Unique Initial Verbs 184 199
Code Segments Mean Std. Dev. Mean Std. Dev.
Lines of Code 29.2 13.7 119.3 75.9
Levenshtein Distance 197.1 260.6 406.6 631.2
Commit Message Mean Std. Dev. Mean Std. Dev.
Tokens 10.1 4.6 23.1 35.2
Table 2. Training dataset statistics for EditPackFT and Commits2023FT
We experiment with two training datasets we gathered: EditPackFT and Commits2023FT, which we describe below. Table 2 presents the statistics for these datasets.
EditPackFT
We created the EditPackFT dataset by further filtering the Python split of the CommitPackFT dataset (Muennighoff et al., 2023), which was used to train OctoCoder. CommitPack is an extensive dataset comprising 4TB of permissively licensed commits from a 2016 GitHub snapshot across various programming languages. CommitPackFT is a subset of CommitPack, filtered for it to be amenable to instruction-tune Code LLMs. The primary criterion for CommitPackFT’s selection involved retaining commits whose messages begin with an imperative verb, mirroring the typical structure of natural language instructions. We apply a series of additional filtering steps, which make the dataset more suitable for code editing. We remove any item that passes any of the following predicates:
@@
    The presence of an empty ‘before’ or ‘after’ code segment, disregarding whitespace.
@@
    No change detected in the ‘before’ and ‘after’ code segments.
@@
    The inclusion of the words TODO, FIXME, or BUG in the ‘after’ code segment, which signals an incomplete commit.
@@
    Incorrect parsing of the ‘after’ code using the Python ast module.
Originally, the dataset contained 56,025 commits, and after applying the filtering steps, we are left with 22,602 commits. As shown by Figure 3(a) and Table 2, the mean number of lines in the ‘before’ and ‘after’ code segments is $29.2$. We further analyzed the original CommitPackFT dataset, ensure that our filtering wasn’t the cause of the short code segments, and find that the mean number of lines is similar, with a mean of $28.8$. This distribution may be suitable for small-scale code editing tasks. We also analyze the distribution of the commit message lengths, and find that the mean token count is $10.1$, which is quite low.
Commits2023
To address the limitations of EditPackFT, we developed the Commits2023FT dataset. This dataset consists of 416,792 Python file changes from commits in permissively licensed GitHub repositories, and is named Commits2023. Our objective is to create a dataset akin to CommitPackFT, but with more recent data and a more diverse example length distribution. We employed the same filters on this dataset as used for EditPackFT, and also applied the initial filters from CommitPackFT, which excludes any commit lacking a message that starts with an imperative verb, or a message that is not within the 10 to 1000 character length range. Additionally, we only retain one file from multi-file commits to avoid exact duplicate commit messages in our dataset (Lee et al., 2022). After this filtering process, we obtain a dataset comprising 24,129 Python file changes. Figure 3(b) presents a broader distribution of the number of lines in the ‘before’ and ‘after’ code segments, with an average of $119.3$. Figure 3 illustrates a sunburst plot of the most frequent initial verbs in the commit messages of Commits2023FT, along with their corresponding root nouns. This set of verbs is slightly more varied than those in EditPackFT, featuring $199$ unique verbs in comparison to $184$. Furthermore, the token count distribution of the commit messages is two times higher and much more varied than that of EditPackFT, with a mean of $23.1$ and a standard deviation of $35.2$.
Ablation Datasets
For ablation analysis, we generated two additional datasets: Commits2023Raw25k and Commits2023FT+EditPackFT. Commits2023Raw25k consists of a random selection of 25,000 commits from Commits2023, we use this dataset to assess the impact of the filtering process on the final dataset. Commits2023FT+EditPackFT represents the combined dataset of Commits2023FT and EditPackFT. In our evaluation, we find that the combination of Commits2023FT and EditPackFT yields the best results by a significant margin (Section 5.2), and thus we train our final models on this dataset. We believe that these results are due to the increased volume of data, and the more varied length distributions.